OpenStudy (moongazer):

Mark invested $4000 in a bank with a 10% monthly interest. How much money will he have after 40 months if the interest covers the previous amount plus interest.

OpenStudy (anonymous):

find the interest for the next 40 months that is 10/100 * 4000 = $400 per month for 40 months = $400 * 40= $16000 Total amount = previous amount + interest = 4000 + 16000 ?

OpenStudy (moongazer):

not that one for example during the first month I got $4200 then for the 2nd month i will have 4200*.05 = 210 4200+210= $4410 I will have $4410 in the 2nd month and so on until I arrive at 40 months

OpenStudy (anonymous):

i think u hve to do it, one by one. since it is a compund interest>

OpenStudy (moongazer):

I got it already :) I found an online calculator that computes things like this. Anyway, thanks for the help. :)

OpenStudy (anonymous):

:)